# Boosting Continual Learning of Vision-Language Models via Mixture-of-Experts Adapters

### v1

@@ E. Effectiveness of Router Selection in MoE-Adapters
   Ours 88.8
59.5
89.1 71.2 65.3 18.2 87.9 44.2 54.6 68.2 64.7(+1.9)
Average    Continual-FT
50.1
56.9
73.5
64.5
45.9
51.2
8.2
81.8
37.9
29.9
38.6
49.0
   LwF [42] 64.1
55.0
79.5
69.2
55.7
58.3
10.8
81.7
41.3
39.2
47.4
54.7
   LwF-VR [15]
63.3
76.9
71.4
79.1
68.9
65.0
13.4
86.0
45.7
36.3
55.3
60.1
   WiSE-FT [67]
59.3
64.7
77.4
70.3
51.3
58.6
10.8
84.2
42.0
38.6
49.1
55.1
@@
   Ours
61.2
87.0 87.3 89.1 79.3 68.5 23.4 89.4 49.9 60.8 68.8 69.5(+1.9)
Last    Continual-FT
35.2
28.6
58.3
51.2
14.0
46.1
5.3
89.5
47.0
52.9
53.6
42.8
   LwF [42]
57.1
40.1
84.1
58.1
50.5
57.6
14.3
87.9
54.7
64.0
47.0
56.8
   LwF-VR [15]
57.3
70.1
72.1
74.6
71.9
65.8
17.4
89.5
60.0
56.0
60.2
63.5
   WiSE-FT [67]
48.1
47.7
66.9
59.8
25.0
56.1
7.4
88.5
52.2
66.8
59.4
51.8
   ZSCL [78] 67.4 82.7
78.7
85.7 81.3 71.2 25.0 92.5 62.0 72.2 74.4 71.8

### v2

@@ E. Effectiveness of Router Selection in MoE-Adapters
     Ours 88.8 59.5 89.1 71.2 65.3 18.2 87.9 44.2 54.6 68.2 64.7(+1.9)
Average      Continual-FT 50.1 56.9 73.5 64.5 45.9 51.2 8.2 81.8 37.9 29.9 38.6 49.0
     LwF [42] 64.1 55.0 79.5 69.2 55.7 58.3 10.8 81.7 41.3 39.2 47.4 54.7
     LwF-VR [15] 63.3 76.9 71.4 79.1 68.9 65.0 13.4 86.0 45.7 36.3 55.3 60.1
     WiSE-FT [67] 59.3 64.7 77.4 70.3 51.3 58.6 10.8 84.2 42.0 38.6 49.1 55.1
@@
     Ours 61.2 87.0 87.3 89.1 79.3 68.5 23.4 89.4 49.9 60.8 68.8 69.5(+1.9)
Last      Continual-FT 35.2 28.6 58.3 51.2 14.0 46.1 5.3 89.5 47.0 52.9 53.6 42.8
     LwF [42] 57.1 40.1 84.1 58.1 50.5 57.6 14.3 87.9 54.7 64.0 47.0 56.8
     LwF-VR [15] 57.3 70.1 72.1 74.6 71.9 65.8 17.4 89.5 60.0 56.0 60.2 63.5
     WiSE-FT [67] 48.1 47.7 66.9 59.8 25.0 56.1 7.4 88.5 52.2 66.8 59.4 51.8
     ZSCL [78] 67.4 82.7 78.7 85.7 81.3 71.2 25.0 92.5 62.0 72.2 74.4 71.8